# It’s Not You, It’s Clipping: A Soft Trust-Region via Probability Smoothing for LLM RL

### v1

@@ 7 Conclusions
We have introduced Probability Smoothing Policy Optimisation (PSPO) as a gradient-preserving alternative to ratio clipping in reinforcement learning for large language models, which mixes the current policy with the behaviour policy before forming the importance ratio. This blend results in a behaviour-anchored soft trust region: it linearly contracts ratios around r=1r{=}1, shrinks TV/KL divergence bounds, and preserves non-zero gradients everywhere.
We empirically evaluate our method by implementing PSPO within GR-PSPO, we consistently outperform clipping GRPO on all datasets, with gains of over $20\%$ on GSM8K on both the 0.5B and 1.5B models. We match the performance of the unclipped, single-pass GRPO, but, importantly, GR-PSPO consistently improves clarity and constraint adherence. Additionally, PSPO is a compute- and memory-neutral, with only a straightforward change to the importance ratio and loss calculations, making it a practical option when multi-epoch updates or mini-batches are used.
Appendix A Full Accuracy Results Table
Model Size Fine-tuning Dataset GSM8K ASDiv SVAMP MATH-500
Smaller Language Models (<1<1B)
GPT-2-Small (Brown et al., ) 0.1B GSM8K & MATH 6.9 - - 5.4*
WizardMath-GPT2-Small (Luo et al., 2025) 0.1B GSM8K & MATH 26.4 - - 12.3*
WizardMath-GPT2-Medium (Luo et al., 2025) 0.3B GSM8K & MATH 38.7 - - 15.6*
GPT-2-Medium (Brown et al., ) 0.3B GSM8K & MATH 11.2 - - 6.2*
GPT-3 (ada) (Ho et al., 2023) 0.3B Distillation of 12 datasets incl. GSM8K & SVAMP 3.1 - 5.0 -
Qwen2.5-0.5B+KD (Zhuang et al., 2025) 0.5B GSM8K 18.7 - - 10.0
Qwen2.5-0.5B+KD (LoRA) (Zhuang et al., 2025) 0.5B GSM8K 9.7 - - 7.6
Qwen2.5-0.5B+SFT (Zhuang et al., 2025) 0.5B GSM8K 21.6 - - 9.2
Qwen2.5-0.5B+SFT (LoRA) (Zhuang et al., 2025) 0.5B GSM8K 2.1 - - 1.2
Qwen2.5-0.5B+RL (Zhuang et al., 2025) 0.5B GSM8K 34.1 - - 23.6
WizardMath-GPT2-Large (Luo et al., 2025) 0.7B GSM8K & MATH 50.1 - - 21.2*
GPT-2-Large (Brown et al., ) 0.7B GSM8K & MATH 13.6 - - 6.4*
FlanT5-Large + Specialised (Fu et al., ) 0.76B GSM8K 20.2 23.8 20.4 -
GPT-2 (Socratic CoT) (Shridhar et al., 2023) 0.77B GSM8K + Socratic CoT 21.1 - - -
GPT-2 (Socratic CoT) (Shridhar et al., 2023) 0.77B SVAMP + Socratic CoT 21.1 - 23.6 -
Larger Language Models (>1>1B)
GPT-3 (babbage) (Ho et al., 2023) 1.3B 12 datasets incl. GSM8K & SVAMP 4.7 - 8.0 -
GPT-2-XL (Brown et al., ) 1.5B GSM8K & MATH 15.4 - - 6.9*
WizardMath-GPT2-XL (Luo et al., 2025) 1.5B GSM8K & MATH 58.9 - - 25.4*
Our Results
Qwen2.5-0.5B 0.5B - 8.3 50.8 26.3 10.2
SFT 0.5B GSM8K 35.7 39.3 14.3
GRPO-clip 0.5B GSM8K 17.6 42.8 34.3 10.3
GRPO-noclip 0.5B GSM8K 40.7 68.3 53.9 17.2
GR-PSPO 0.5B GSM8K 39.7 68.1 53.2 16.8
Qwen2.5-1.5B 1.5B - 1.6 2.4 1.7 5.1
GRPO-clip 1.5B GSM8K 37.8 70.9 58.4 14.9
GRPO-noclip 1.5B GSM8K 57.9 80.4 74.9 25.2
GR-PSPO 1.5B GSM8K 59.4 77.7 70.3 25.6
Table 5: Performance of similar-size models from the literature, compared with our results. We report our model accuracy from the zero-shot Top-1 accuracy.
*These models were evaluated against MATH rather than MATH-500. Where MATH-500 is a subset of MATH, we still compare the results here.
Table 5 contains all of our results alongside all published results we compared with in our results section.
Appendix B LLM-as-Judge Setup
We use the following rubric:
You are an evaluator. Judge ONLY the reasoning quality (not final answer).
Score four criteria from 0 (bad) to 5 (excellent):
1) Uses given constraints/numbers
2) Logical flow (no leaps/contradictions)
3) Mathematical soundness of steps
4) Clarity/conciseness (limited repetition/hedging)
Return strict JSON: ”score”:”constraints”:X,”logic”:Y,”math”:Z,”clarity”:W,”overall”:O,”rationale”:”…”.
Overall is the rounded mean of the four sub-scores.
We use the following system prompt:
You are a fair, strict, and concise evaluator. Output ONLY JSON.
We use the following user prompt:
Problem: prompt
Model reasoning: completion
Provide your JSON now.
We used llama-3.3-70b-versatile as our judge model, accessing it via an API service. We averaged scores per dataset per model across all responses and seeds. We used the $T=0$ answers.

### v2

@@ 6 Conclusions
We introduced Probability Smoothing Policy Optimisation (PSPO) as a gradient-preserving alternative to ratio clipping in reinforcement learning for large language models, which mixes the current policy with the behaviour policy before forming the importance ratio. This blend results in a behaviour-anchored soft trust region that linearly contracts ratios around r=1r{=}1, shrinks TV/KL divergence bounds (Lemma 2.1, Corollary 2.2), and preserves non-zero gradients everywhere (Proposition 2.3).
We empirically evaluate our method by implementing PSPO within GRPO as GR-PSPO, our method demonstrates clear advantages in settings where stability matters most. On Qwen2-Math-1.5B, a model with prior supervised mathematical training, GR-PSPO outperforms baselines across all four benchmarks; giving results similar to Llama 3 8B+TOPR Roux and Bellemare (2025). On base Qwen2.5-1.5B, GR-PSPO achieves the strongest MATH performance, exceeding GRPO-clip by 4.1pp. These results support our theoretical motivation that when there is existing capability to preserve, PSPO’s behaviour-anchored smoothing prevents the destabilising updates more effectively than clipping.
Additionally, PSPO is a drop-in replacement for clipping. PSPO requires tuning of only one hyperparameter ($\alpha$, replacing $\varepsilon$), and adds no computational or memory overhead. These properties make it particularly suited to settings where data reuse is necessary and preserving existing model capabilities is critical, the common case in post-SFT reinforcement learning pipelines.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Training Details
The training for each method was run on 1 GPU. We used python version 3.10.18, and trl=0.25.1. We save checkpoint models every 1000 steps, and return the best model checkpoint at the end of training.
We used Nvidia GPUs (A100) and AMD GPUs (MI300X) for training, using only 1 GPU at a time. It took approximately 18 hours per training run for the 150M tokens, and around 10 hours for the 75M token training runs.
A.1 Tuning Method
Each method was independently tuned using a grid search over learning rate and any parameter related to the trust region (i.e. clipping range, smoothing alpha value, tau values).
We ran the initial grid search using seed 42, for 500 max global steps. The top 5 configurations were then run again on seeds 43 and 44. The best average validation reward from this was selected for the full training runs. These seeds are different to those used during training (seeds 0, 1 and 2 were used in training).
We used the learning rates =[5e07,1e06,5e06]=[5e-07,1e-06,5e-06], and the method specific values used are shown in Table 4.
Table 4: Hyperparameter grid search values for method specific parameters.
Method Trust Region Parameter Grid Search Values Notes
GRPO-clip clipping range, $\varepsilon$ [0.1, 0.2, 0.3, 0.4]
GRPO-noclip N/A N/A
P3O temperature, $\tau$ [1.0, 2.0, 4.0, 10.0] Chen et al. (2023) suggested $\tau=4.0$
SAPO temperature for positive advantages, $\tau_{positive}$ [1.0, 2.0, 3.0] Gao et al. (2025) suggested that $\tau_{negative}>\tau_{positive}$ gave the most stable training, so we only considered configurations where $\tau_{negative}\geq\tau_{positive}$.
temperature for negative advantages, $\tau_{negative}$ [1.0, 2.0, 3.0]
TOPR truncation limit for positive advantages, $b^{+}$ [1.0, 2.0, 10.0] Roux and Bellemare (2025) suggested $a^{-}=0$ and $a^{+}=1$; we searched across $b^{+}$ and $b^{-}$ values (paper uses $b^{+}=b^{-}=1$).
truncation limit for negative advantages, $b^{-}$ [1.0, 2.0, 10.0]
GR-PSPO smoothing alpha, $\alpha$ [0.1, 0.2, 0.3, 0.4]
A.2 Hyperparameter Values
A.2.1 Qwen2.5-0.5B and Qwen2.5-1.5B
For the Qwen2.5-0.5B and -1.5B models, we used the shared parameters shown in Table 5.
Table 5: Shared training hyperparameters for the 0.5B and 1.5B experiments.
Parameter Value
Effective batch size 128 (16 per device $\times$ 8 accumulation steps)
Generations per prompt 4
KL penalty ($\beta$) 0.0
Max completion length 128 tokens
Sampling temperature 0.6
Top-pp 0.85
Warmup steps 125
Precision bfloat16
The method specific parameters are shown in Table 6.
Table 6: Hyperparameters for each trust region method for Qwen2.5-0.5B and -1.5B. All methods use the DAPO loss unless otherwise noted.
Method Iterations Learning Rate Trust Region Parameters
GRPO-noclip 1 1e-6
GRPO-clip 2 1e-6 $\varepsilon=0.4$
GR-PSPO 2 1e-6 $\alpha=0.1$
P3O 2 1e-6 $\tau=4.0$
SAPO 2 1e-6 $\tau_{+}=1.0$, $\tau_{-}=3.0$
TOPR 2 1e-6 $b^{+}=1.0$, $b^{-}=1.0$
Uses GRPO-style loss to match existing implementations.
A.2.2 Qwen2-Math-1.5B
For the Qwen2-Math-1.5B model, we used the shared parameters shown in Table 5.
The method specific parameters are shown in Table 7.
Table 7: Hyperparameters for each trust region method for Qwen2-Math-1.5B. All methods use the DAPO loss unless otherwise noted.
Method Iterations Learning Rate Trust Region Parameters
GRPO-noclip 1 5e-6
GRPO-clip 2 5e-6 $\varepsilon=0.2$
GR-PSPO 2 1e-6 $\alpha=0.4$
P3O 2 1e-6 $\tau=4.0$
SAPO 2 1e-6 $\tau_{+}=1.0$, $\tau_{-}=3.0$
TOPR 2 1e-6 $b^{+}=1.0$, $b^{-}=1.0$
Uses GRPO-style loss to match existing implementations.
Appendix B Sensitivity to Smoothing Alpha
We ablate the smoothing parameter $\alpha$ on Qwen2-Math-1.5B with a learning rate of 1e-6 for 250 global steps. Performance peaks around $\alpha$=0.4, with degradation at both extremes, which could support our theory that insufficient smoothing ($\alpha$=0.1) provides weak regularisation, while excessive smoothing ($\alpha$=0.5) over-constrains updates. Table 8 shows the mean validation reward at the end of the 250 steps across different values of $\alpha$.
Table 8: Small ablation across smoothing alpha, $\alpha$, values on Qwen2-Math-1.5B.
$\alpha$ Mean Validation Reward
0.1 0.170
0.2 0.276
0.4 0.319
0.5 0.240